OpenStudy (anonymous):

Convert. If necessary, round to the nearest tenth (Recall: 1 lb. = 16 oz. and and 1 g .035 ounces) 500 g. lb. a. 5.5 c. 1.1 b. 1.5 d. 5.1

OpenStudy (anonymous):

What is your question asking \[500g.ib.?\] grams and pounds, can you please specify

OpenStudy (anonymous):

Is that suppose to be 500 g and 1 pound?

OpenStudy (anonymous):

Well to convert the 500 lbs into ounces notice how they gave you "recall 1g = 0.035 ounces" So all you have to do to convert grams into ounces is multiply. So we have \[500 g \times \frac{ 0.035 oz }{ 1 g } = 17.5 oz \] the ratio \[\frac{ 0.035oz }{ 1g }\] implies in 1 gram there are 0.035 ounces. And we multiply over 1g when we convert to cancel out the grams. Does that make sense? And here this will help you round to tenths, 

OpenStudy (anonymous):

Sorry that should be written as to convert 500g to ounces not pounds :)

OpenStudy (anonymous):

Now how many pounds do you need to convert? Notice that you would do the same thing as before except now we have 1 lb = 16 oz

OpenStudy (anonymous):

Your image just shows a grey box

OpenStudy (anonymous):

Oh I see! They want you to convert 500g into pounds!

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

Ok so we have converted grams into ounces, so we have 500g = 17.5 oz, now we can convert these 17.5 oz into lbs, yeah it seems weird because we're going backwards, but that's how they want it. Ok notice how it says \[1 lb = 16 oz\] since we're going the other way, all we have to do is divide then \[17.5 oz \times \frac{ 1 lb }{ 16 oz } \approx ...lb\]

OpenStudy (anonymous):

So how many pounds is that?

OpenStudy (anonymous):

Let me know if you don't understand something, I can try explaining it better :)

OpenStudy (anonymous):

so just multiply 17.5oz * .0625 and it wil give me the answer

OpenStudy (anonymous):

Yes, exactly!

OpenStudy (anonymous):

1.093?

OpenStudy (anonymous):

Yes, now round that to the tenth

OpenStudy (anonymous):



OpenStudy (anonymous):

1.1

OpenStudy (anonymous):

Convert both 16.1 km/h and 880 ft/min to meters per second and compare.